# A virtual Clebsch-Gordan formula for $\operatorname{SL}(2,\mathbb{Z})$

### v5

@@ Abstract.
If $k$ is a field of characteristic zero and if $\operatorname{SL}(n,k)$ is the group of $n\times n$ matrices with coefficients in $k$ and determinant one, it follows $\operatorname{SL}(n,k)$ is ”semi simple” in the sense that any finite dimensional representation decompose into a finite direct sum of irreducible representations. Moreover any surjective map $p:V\rightarrow W$ of representations has a section $s:W\rightarrow V$ with $s\circ p=Id$ where $s$ is a map of representations. We prove that this property no longer holds for the group scheme $G:=\operatorname{SL}(2,\mathbb{Z})$ by giving explicit examples of a surjective map of comodules that does not have a section. Hence $G$ is not ”completely reducible” in this sense. We also prove this result for $\operatorname{SL}(2,\mathbb{Z}[\frac{1}{m}])$ and for $\operatorname{SL}(2,\mathbb{Z}_{(p)})$ where $\mathbb{Z}_{(p)}$ is the localization of $\mathbb{Z}$ at the prime ideal $(p)$. If $V:=\mathbb{Z}\{e_{1},e_{2}\}$ is the standard comodule we may form the symmetric product $\operatorname{S}^{n}(V)$, and over a field of characteristic zero it follows this construction gives all finite dimensional irreducible modules. Over the integers it follows $\operatorname{S}^{n}(V)$ is no longer irreducible as a comodule, hence there is no classification of $\operatorname{SL}(2,\mathbb{Z})$-modules as in the case of a field of characteristic zero. There is a version of the Clebsch-Gordan formula in the following sense: There is for any pair of integers $n,m$ with $1\leq n\leq m$ a finite filtration $F_{i}\subseteq\operatorname{S}^{n}(V)\otimes\operatorname{S}^{m}(V)$ of comodules, with $F_{i}/F_{i+1}\cong\operatorname{S}^{n+m-2i}(V)$. Hence in the grothendieck group we get an equality of classes - a virtual Clenbsch-Gordan formula:
@@
Hence the Clebsch-Gordan formula holds in the grothendieck group of finite rank comodules on $\operatorname{SL}(2,\mathbb{Z})$. The filtration $F_{i}$ pulls back to a filtration $F_{i}\otimes_{\mathbb{Z}}B$ of $\operatorname{S}^{n}(V_{B})\otimes\operatorname{S}^{m}(V_{B})$ for any commutative ring $B$, and it gives a virtual Clebsch-Gordan formula in the grothendieck group of ”finite rank” comodules on $\operatorname{SL}(2,B)$ for any such $B$. Here $V_{B}:=V\otimes_{\mathbb{Z}}B$.
We introduce the notion ”good filtration” for a torsion free comodule on $\operatorname{SL}(2,\mathbb{Z})$ and give non trivial examples of torsion free modules equipped with a good filtration.
@@ 1. Introduction
The aim of the paper is to study torsion free finite rank comodules on the group scheme $\operatorname{SL}(2,A)$ where $A$ is a Dedekind domain. If $A$ is a field of characteristic zero there is a complete classification of finite dimensional irreducbible comodules. For any such comodule $W$ there is a direct sum decomposition $W\cong\oplus_{i=1}^{n}\operatorname{S}^{d_{i}}(V)$ where $V$ is the ”standard comodule” and $\operatorname{S}^{d}(V)$ is the d’th symmetric power of $V$. Moreover for any integer $d\geq 0$ it follows the comodule $\operatorname{S}^{d}(V)$ is an irreducible comodule. These properties no longer hold for a general Dedekind domain such as $\mathbb{Z},\mathbb{Z}_{(p)}$ or $\mathbb{Z}[\frac{1}{m}]$. For a torsion free finite rank comodule $W$ on $\operatorname{SL}(2,A)$ there is no direct sum decomposition into symmetric powers of the standard comodule $V$ (see Theorem 3.4). Moreover the symmetric powers $\operatorname{S}^{d}(V)$ are no longer irreducible comodules on $\operatorname{SL}(2,A)$ (see Example 3.7).
@@
of comodules on $\operatorname{SL}(2,A)$. This formula gives many filtrations of the tensor product. One of them is the following filtration: Let $F(k)_{i}:=\oplus_{j=0}^{n}\operatorname{S}^{n+m-2j}(V)$ It follows we get a filtration - the classical Clebsch-Gordan filtration -
@@
This filtration is a filtration of $G$-comodules with $F_{i}/F_{i+1}\cong\operatorname{S}^{n+m-2i}(V)$ for all $i=0,..,n$ (see Corollary 2.9). The filtration $F_{i}$ pulls back to the ”classical Clebsch-Gordan filtration” filtration in the case when $k$ is a field of characteristic zero. It follows the universal Clebsch-Gordan filtration is a good filtration of comodules of $\operatorname{S}^{n}(V)\otimes\operatorname{S}^{m}(V)$ for the group scheme $\operatorname{SL}(2,A)$ over any commutative ring $A$. Hence the Clebsch-Gordan decomposition does not generalize to an arbitrary commutative ring but the classical Clebsch-Gordan filtration does.
We study the ”complete reducibility property” for group schemes $\operatorname{SL}(2,A)$ where $A:=\mathbb{Z},\mathbb{Z}_{(p)},\mathbb{Z}[\frac{1}{m}]$, and find that these group schemes are not ”completely reducible” in the sense that there are surjections of comodules $\pi:V\rightarrow W$ that do not have a section (see Theorem 3.4). Moreover: The comodule $\operatorname{S}^{n}(V)$ is no longer irreducible when $A$ is one of the rings $\mathbb{Z},\mathbb{Z}_{(p)}$ or $\mathbb{Z}[\frac{1}{m}]$. Hence there is no classification of irreducible $\operatorname{SL}(2,A)$-comodules in terms of direct sums of the symmetric powers $\operatorname{S}^{n}(V)$ of the standard module $V$.
We give explicit examples of comodules $W,W^{*}$ on $\operatorname{SL}(2,\mathbb{Z})$ with the property that $W$ and $W^{*}$ have the same weight space decomposition but where $W$ and $W^{*}$ are non isomorphic as comodules. There is an isomorphism $W\otimes_{\mathbb{Z}}\mathbb{Z}[\frac{1}{2}]\cong W^{*}\otimes_{\mathbb{Z}}\mathbb{Z}[\frac{1}{2}]$ as comodules on $\operatorname{SL}(2,\mathbb{Z}[\frac{1}{2}])$. Hence the modules $W,W^{*}$ are not classified by their weights and weight spaces (see Theorem 5.4). We also give examples of comodules with no good filtration (see Example 5.3).
Note: In a paper from 2020 (see [9]) I initiated a study of flag bundles $\mathbb{F}(E)$ where $E$ is a finite rank vector bundle on $S:=\operatorname{Spec}(A)$ and where $A$ is a Dedekind domain. If $P\subseteq\operatorname{SL}(n,A)$ is a parabolic subgroup scheme of $\operatorname{SL}(n,A)$, it follows the quotient $\operatorname{SL}(n,A)/P$ is such a flag bundle. Any representation $(V,\rho)$ of $P$ with the property that $V$ is a torsion free and finite rank projective $A$-module, gives rise to a finite rank vector bundle $E(\rho)$ on $\operatorname{SL}(n,A)/P$ and for this reason one wants to be able to classify such representations. It is hoped the constructions introduced in this paper may have applications in the study of such flag bundles and their vector bundles.
@@ 2. Symmetric powers and the universal Clebsch-Gordan filtration
In this section we wil study representations of the group scheme $\operatorname{SL}(2,A)$ where $A$ is any commutative ring and $\operatorname{SL}(2,\mathbb{Z})$ where $\mathbb{Z}$ is the ring of integers. We will us the language of ”affine group schemes” introduced in the books of Demazure, Gabriel and Jantzen (see [3] and [8]). If $G:=\operatorname{Spec}(R)$ is an affine group scheme with Hopf algebra $R$ we will let $h_{R}$ denote the ”functor of points” of the group scheme $G$.
Let in the following $A$ be a fixed commutative unital ring and let $V:=A\{e_{1},e_{2}\}$ be the free $A$-module on the elements $e_{1},e_{2}$. Let $x_{11},x_{12},x_{21},x_{22}$ be independent variables over $A$ and let $D:=x_{11}x_{22}-x_{12}x_{21}$ Let $R:=A[x_{ij}]/(D-1)$ and let $\operatorname{SL}(2,A):=\operatorname{Spec}(R)$. It follows the functor of points
@@
Hence $(V,\Delta_{V})$ is a comodule on $\operatorname{SL}(2,A)$.
@@
Let $(V,\Delta_{V})$ be the standard comodule on $\operatorname{SL}(2,A)$.
@@
The structure $\Delta_{V}$ as comodule on $V$ corresponds to ”matrix multiplication from the right”. Let $v\in V$ be represented by a pair $[a_{1},a_{2}]$ with $a_{i}\in A$, and choose a $A$-valued point $M\in\operatorname{SL}(2,A)(A)$, where $M$ is a $2\times 2$-matrix with coefficients in $A$ and determinant one. The matrix $M$ acts on the vector $[a_{1},a_{2}]$ using matrix multiplication from the right. There is another obvious comodule structure
@@
defined by $f(u\otimes v):=u\wedge v$ is a nondegenerate pairing and the fact that $\wedge^{2}V\cong T$ is the trivial comodule. This result is by the above argument true for $\operatorname{SL}(2,A)$ over any commutative ring $A$.
@@
Since $G:=\operatorname{SL}(2,\mathbb{Z})$ is a group scheme over $\mathbb{Z}$ it follows its algebra of distributions $\operatorname{Dist}(G)$ acts on $V$ and $\operatorname{S}^{d}(V)$. There is an inclusion $\operatorname{Lie}(G)\subseteq\operatorname{Dist}(G)$ and $\operatorname{Lie}(G)\cong\mathfrak{sl}(2,\mathbb{Z})$ is the Lie algebra of $2\times 2$-matrices with coefficients in $\mathbb{Z}$ and trace equal to zero. It follows $\operatorname{S}^{d}(V)$ is a module on $\operatorname{Lie}(G)$ and it has by [3] a weight space decomposition for the action of $H$. We let $\operatorname{Lie}(G):=\mathbb{Z}\{x,H,y\}$ be the ”standard generators” for $\operatorname{Lie}(G)$. The elements $x,H,y$ are defined as follows:
@@
Let $1\leq n\leq m$ be a pair of integers. There is an exact sequence of $\operatorname{SL}(2,\mathbb{Z})$-comodules
@@
Let $1\leq n\leq m$ be a pair of integers. There is a filtration of $\operatorname{SL}(2,\mathbb{Z})$-comodules
@@
Let $1\leq n\leq m$ be a pair of integers. The filtration of $\operatorname{SL}(2,\mathbb{Z})$-comodules
@@
(The virtual Clebsch-Gordan formula) Let $1\leq n\leq m$ be a pair of integers. Let $\operatorname{K}(\operatorname{SL}(2,\mathbb{Z}))$ be the grothendieck group of finite rank comodules on $\operatorname{SL}(2,\mathbb{Z})$. There is an equality
@@
in $\operatorname{K}(\operatorname{SL}(2,\mathbb{Z}))$.
@@
Let $S.=\operatorname{Spec}(\mathbb{Z})$ and let $\pi:\operatorname{SL}(2,\mathbb{Z}):=G\rightarrow S$ be the canonical map. Let $k$ be the field of rational numbers and let $T:=\operatorname{Spec}(k)$. It follows there is an isomorphism $G_{k}:=\pi^{-1}(T)\cong\operatorname{SL}(2,k)$ and we may view $G_{k}$ as the generic fiber of the map $\pi$. The group scheme $G_{k}$ is completely reducible and we may pull back $\operatorname{S}^{n}(V)\otimes\operatorname{S}^{m}(V)$ to $G_{k}$ to get the comodule $\operatorname{S}^{n}(V_{k})\otimes_{k}\operatorname{S}^{m}(V_{k})$ where $V_{k}:=V\otimes_{\mathbb{Z}}k$. The Clebsch-Gordan formula holds for $\operatorname{S}^{n}(V_{k})\otimes\operatorname{S}^{m}(V_{k})$: Let $1\leq n\leq m$ be a pair of integers. There is an isomorphism of $G_{k}$-comodules
@@
for all $i=0,..,n$. Hence the Clebsch-Gordan formula does not generalize to $\operatorname{SL}(2,\mathbb{Z})$ but the filtration does. There are many different filtrations of the comodule $\operatorname{S}^{n}(V_{k})\otimes\operatorname{S}^{m}(V_{k})$. The filtration $F(k)_{i}$ gives an ordering of the direct summands $\operatorname{S}^{n+m-2i}(V_{k})$ of $\operatorname{S}^{n}(V_{k})\otimes\operatorname{S}^{m}(V_{k})$.
@@
3. On the ”complete reduciblilty” property for $\operatorname{SL}(2,\mathbb{Z})$
Let in this section $p\in\mathbb{Z}$ be a prime number and let $\mathbb{Z}_{(p)}$ be the localization of $\mathbb{Z}$ at $p$. Let $m$ be an integer and let $\mathbb{Z}[\frac{1}{m}]$ be the localization of $\mathbb{Z}$ at the set $S:=\{m^{i}\}$. We prove that the group schemes $\operatorname{SL}(2,\mathbb{Z}),\operatorname{SL}(2,\mathbb{Z}_{(p)})$ and $\operatorname{SL}(2,\mathbb{Z}[\frac{1}{m}])$ do not satisfy the ”complete reducibility” property. Over a field of $k$ characteristic zero, any finite dimensional comodule on $\operatorname{SL}(2,k)$ satisfies this property. In order to study torsion free comodules on $\operatorname{SL}(2,\mathbb{Z})$ we introduce the notion ”good filtration” and discuss the problem of classifying torsion free finite rank comodules with a good filtration. We give explicit non trivial examples of torsion free finite rank module with a good filtration, hence many nontrivial torsion free comodules has such a filtration, and the notion may therefore have general interest in the study of comodules on a reductive group scheme over a Dedekind domain. One may define the notion ”good filtration” for torsion free comodules on $\operatorname{SL}(n,A)$ for any Dedekind domain $A$.

### v6

# A virtual Clebsch-Gordan formula for $\operatorname{SL}_{2,A}$

@@ Abstract.
If $k$ is a field of characteristic zero and if $\operatorname{SL}(n,k)$ is the group of $n\times n$ matrices with coefficients in $k$ and determinant one, it follows $\operatorname{SL}(n,k)$ is ”semi simple” in the sense that any finite dimensional representation decompose into a finite direct sum of irreducible representations. Moreover any surjective map $p:V\rightarrow W$ of representations has a section $s:W\rightarrow V$ with $s\circ p=Id$ where $s$ is a map of representations. We prove that this property no longer holds for the group scheme $G:=\operatorname{SL}_{2,\mathbb{Z}}$ by giving explicit examples of a surjective map of comodules that does not have a section. Hence $G$ is not ”completely reducible” in this sense. We also prove this result for $\operatorname{SL}_{2,\mathbb{Z}[\frac{1}{m}]}$ and for $\operatorname{SL}_{2,\mathbb{Z}_{(p)}}$ where $\mathbb{Z}_{(p)}$ is the localization of $\mathbb{Z}$ at the prime ideal $(p)$. If $V:=\mathbb{Z}\{e_{1},e_{2}\}$ is the standard comodule we may form the symmetric product $\operatorname{S}^{n}(V)$, and over a field of characteristic zero it follows this construction gives all finite dimensional irreducible modules. Over the integers it follows $\operatorname{S}^{n}(V)$ is no longer irreducible as a comodule, hence there is no classification of $\operatorname{SL}_{2,\mathbb{Z}}$-modules as in the case of a field of characteristic zero. There is a version of the Clebsch-Gordan formula in the following sense: There is for any pair of integers $n,m$ with $1\leq n\leq m$ a finite filtration $F_{i}\subseteq\operatorname{S}^{n}(V)\otimes\operatorname{S}^{m}(V)$ of comodules, with $F_{i}/F_{i+1}\cong\operatorname{S}^{n+m-2i}(V)$. Hence in the grothendieck group we get an equality of classes - a virtual Clenbsch-Gordan formula:
@@
Hence the Clebsch-Gordan formula holds in the grothendieck group of finite rank comodules on $\operatorname{SL}_{2,\mathbb{Z}}$. The filtration $F_{i}$ pulls back to a filtration $F_{i}\otimes_{\mathbb{Z}}A$ of $\operatorname{S}^{n}(V_{A})\otimes\operatorname{S}^{m}(V_{A})$ for any commutative ring $A$, and it gives a virtual Clebsch-Gordan formula in the grothendieck group of ”finite rank” comodules on $\operatorname{SL}_{2,A}$ for any such $A$. Here $V_{A}:=V\otimes_{\mathbb{Z}}A$.
We introduce the notion ”good filtration” for a torsion free comodule on $\operatorname{SL}_{2,\mathbb{Z}}$ and give non trivial examples of torsion free modules equipped with a good filtration.
@@ 1. Introduction
The aim of the paper is to study torsion free finite rank comodules on the group scheme $\operatorname{SL}_{2,A}$ where $A$ is a Dedekind domain. If $A$ is a field of characteristic zero there is a complete classification of finite dimensional irreducbible comodules. For any such comodule $W$ there is a direct sum decomposition $W\cong\oplus_{i=1}^{n}\operatorname{S}^{d_{i}}(V)$ where $V$ is the ”standard comodule” and $\operatorname{S}^{d}(V)$ is the d’th symmetric power of $V$. Moreover for any integer $d\geq 0$ it follows the comodule $\operatorname{S}^{d}(V)$ is an irreducible comodule. These properties no longer hold for a general Dedekind domain such as $\mathbb{Z},\mathbb{Z}_{(p)}$ or $\mathbb{Z}[\frac{1}{m}]$. For a torsion free finite rank comodule $W$ on $\operatorname{SL}_{2,A}$ there is no direct sum decomposition into symmetric powers of the standard comodule $V$ (see Theorem 3.4). Moreover the symmetric powers $\operatorname{S}^{d}(V)$ are no longer irreducible comodules on $\operatorname{SL}_{2,A}$ (see Example 3.7).
@@
of comodules on $\operatorname{SL}_{2,A}$. This formula gives many filtrations of the tensor product. One of them is the following filtration: Let $F(k)_{i}:=\oplus_{j=0}^{n}\operatorname{S}^{n+m-2j}(V)$ It follows we get a filtration - the classical Clebsch-Gordan filtration -
@@
This filtration is a filtration of $G$-comodules with $F_{i}/F_{i+1}\cong\operatorname{S}^{n+m-2i}(V)$ for all $i=0,..,n$ (see Corollary 2.9). The filtration $F_{i}$ pulls back to the ”classical Clebsch-Gordan filtration” filtration in the case when $k$ is a field of characteristic zero. It follows the universal Clebsch-Gordan filtration is a good filtration of comodules of $\operatorname{S}^{n}(V)\otimes\operatorname{S}^{m}(V)$ for the group scheme $\operatorname{SL}_{2,A}$ over any commutative ring $A$. Hence the Clebsch-Gordan decomposition does not generalize to an arbitrary commutative ring but the classical Clebsch-Gordan filtration does.
We study the ”complete reducibility property” for group schemes $\operatorname{SL}_{2,A}$ where $A:=\mathbb{Z},\mathbb{Z}_{(p)},\mathbb{Z}[\frac{1}{m}]$, and find that these group schemes are not ”completely reducible” in the sense that there are surjections of comodules $\pi:V\rightarrow W$ that do not have a section (see Theorem 3.4). Moreover: The comodule $\operatorname{S}^{n}(V)$ is no longer irreducible when $A$ is one of the rings $\mathbb{Z},\mathbb{Z}_{(p)}$ or $\mathbb{Z}[\frac{1}{m}]$. Hence there is no classification of irreducible $\operatorname{SL}_{2,A}$-comodules in terms of direct sums of the symmetric powers $\operatorname{S}^{n}(V)$ of the standard module $V$.
We give explicit examples of comodules $W,W^{*}$ on $\operatorname{SL}_{2,\mathbb{Z}}$ with the property that $W$ and $W^{*}$ have the same weight space decomposition but where $W$ and $W^{*}$ are non isomorphic as comodules. There is an isomorphism $W\otimes_{\mathbb{Z}}\mathbb{Z}[\frac{1}{2}]\cong W^{*}\otimes_{\mathbb{Z}}\mathbb{Z}[\frac{1}{2}]$ as comodules on $\operatorname{SL}_{2,\mathbb{Z}[\frac{1}{2}]}$. Hence the modules $W,W^{*}$ are not classified by their weights and weight spaces (see Theorem 5.4). We also give examples of comodules with no good filtration (see Example 5.3).
Note: In a paper from 2020 (see [9]) I initiated a study of flag bundles $\mathbb{F}(E)$ where $E$ is a finite rank vector bundle on $S:=\operatorname{Spec}(A)$ and where $A$ is a Dedekind domain. If $P\subseteq\operatorname{SL}_{n,A}$ is a parabolic subgroup scheme of $\operatorname{SL}_{n,A}$, it follows the quotient $\operatorname{SL}_{n,A}/P$ is such a flag bundle. Any representation $(V,\rho)$ of $P$ with the property that $V$ is a torsion free and finite rank projective $A$-module, gives rise to a finite rank vector bundle $E(\rho)$ on $\operatorname{SL}_{n,A}/P$ and for this reason one wants to be able to classify such representations. It is hoped the constructions introduced in this paper may have applications in the study of such flag bundles and their vector bundles.
@@ 2. Symmetric powers and the universal Clebsch-Gordan filtration
In this section we wil study representations of the group scheme $\operatorname{SL}_{2,A}$ where $A$ is any commutative ring and $\operatorname{SL}_{2,\mathbb{Z}}$ where $\mathbb{Z}$ is the ring of integers. We will us the language of ”affine group schemes” introduced in the books of Demazure, Gabriel and Jantzen (see [3] and [8]). If $G:=\operatorname{Spec}(R)$ is an affine group scheme with Hopf algebra $R$ we will let $h_{R}$ denote the ”functor of points” of the group scheme $G$.
Let in the following $A$ be a fixed commutative unital ring and let $V:=A\{e_{1},e_{2}\}$ be the free $A$-module on the elements $e_{1},e_{2}$. Let $x_{11},x_{12},x_{21},x_{22}$ be independent variables over $A$ and let $D:=x_{11}x_{22}-x_{12}x_{21}$ Let $R:=A[x_{ij}]/(D-1)$ and let $\operatorname{SL}_{2,A}:=\operatorname{Spec}(R)$. It follows the functor of points
@@
Hence $(V,\Delta_{V})$ is a comodule on $\operatorname{SL}_{2,A}$.
@@
Let $(V,\Delta_{V})$ be the standard comodule on $\operatorname{SL}_{2,A}$.
@@
The structure $\Delta_{V}$ as comodule on $V$ corresponds to ”matrix multiplication from the right”. Let $v\in V$ be represented by a pair $[a_{1},a_{2}]$ with $a_{i}\in A$, and choose a $A$-valued point $M\in\operatorname{SL}_{2,A}(A)$, where $M$ is a $2\times 2$-matrix with coefficients in $A$ and determinant one. The matrix $M$ acts on the vector $[a_{1},a_{2}]$ using matrix multiplication from the right. There is another obvious comodule structure
@@
defined by $f(u\otimes v):=u\wedge v$ is a nondegenerate pairing and the fact that $\wedge^{2}V\cong T$ is the trivial comodule. This result is by the above argument true for $\operatorname{SL}_{2,A}$ over any commutative ring $A$.
@@
Since $G:=\operatorname{SL}_{2,\mathbb{Z}}$ is a group scheme over $\mathbb{Z}$ it follows its algebra of distributions $\operatorname{Dist}(G)$ acts on $V$ and $\operatorname{S}^{d}(V)$. There is an inclusion $\operatorname{Lie}(G)\subseteq\operatorname{Dist}(G)$ and $\operatorname{Lie}(G)\cong\mathfrak{sl}(2,\mathbb{Z})$ is the Lie algebra of $2\times 2$-matrices with coefficients in $\mathbb{Z}$ and trace equal to zero. It follows $\operatorname{S}^{d}(V)$ is a module on $\operatorname{Lie}(G)$ and it has by [3] a weight space decomposition for the action of $H$. We let $\operatorname{Lie}(G):=\mathbb{Z}\{x,H,y\}$ be the ”standard generators” for $\operatorname{Lie}(G)$. The elements $x,H,y$ are defined as follows:
@@
Let $1\leq n\leq m$ be a pair of integers. There is an exact sequence of $\operatorname{SL}_{2,\mathbb{Z}}$-comodules
@@
Let $1\leq n\leq m$ be a pair of integers. There is a filtration of $\operatorname{SL}_{2,\mathbb{Z}}$-comodules
@@
Let $1\leq n\leq m$ be a pair of integers. The filtration of $\operatorname{SL}_{2,\mathbb{Z}}$-comodules
@@
(The virtual Clebsch-Gordan formula) Let $1\leq n\leq m$ be a pair of integers. Let $\operatorname{K}(\operatorname{SL}_{2,\mathbb{Z}})$ be the grothendieck group of finite rank comodules on $\operatorname{SL}_{2,\mathbb{Z}}$. There is an equality
@@
in $\operatorname{K}(\operatorname{SL}_{2,\mathbb{Z}})$.
@@
Let $S.=\operatorname{Spec}(\mathbb{Z})$ and let $\pi:\operatorname{SL}_{2,\mathbb{Z}}:=G\rightarrow S$ be the canonical map. Let $k$ be the field of rational numbers and let $T:=\operatorname{Spec}(k)$. It follows there is an isomorphism $G_{k}:=\pi^{-1}(T)\cong\operatorname{SL}(2,k)$ and we may view $G_{k}$ as the generic fiber of the map $\pi$. The group scheme $G_{k}$ is completely reducible and we may pull back $\operatorname{S}^{n}(V)\otimes\operatorname{S}^{m}(V)$ to $G_{k}$ to get the comodule $\operatorname{S}^{n}(V_{k})\otimes_{k}\operatorname{S}^{m}(V_{k})$ where $V_{k}:=V\otimes_{\mathbb{Z}}k$. The Clebsch-Gordan formula holds for $\operatorname{S}^{n}(V_{k})\otimes\operatorname{S}^{m}(V_{k})$: Let $1\leq n\leq m$ be a pair of integers. There is an isomorphism of $G_{k}$-comodules
@@
for all $i=0,..,n$. Hence the Clebsch-Gordan formula does not generalize to $\operatorname{SL}_{2,\mathbb{Z}}$ but the filtration does. There are many different filtrations of the comodule $\operatorname{S}^{n}(V_{k})\otimes\operatorname{S}^{m}(V_{k})$. The filtration $F(k)_{i}$ gives an ordering of the direct summands $\operatorname{S}^{n+m-2i}(V_{k})$ of $\operatorname{S}^{n}(V_{k})\otimes\operatorname{S}^{m}(V_{k})$.
@@
3. On the ”complete reduciblilty” property for $\operatorname{SL}_{2,\mathbb{Z}}$
Let in this section $p\in\mathbb{Z}$ be a prime number and let $\mathbb{Z}_{(p)}$ be the localization of $\mathbb{Z}$ at $p$. Let $m$ be an integer and let $\mathbb{Z}[\frac{1}{m}]$ be the localization of $\mathbb{Z}$ at the set $S:=\{m^{i}\}$. We prove that the group schemes $\operatorname{SL}_{2,\mathbb{Z}},\operatorname{SL}_{2,\mathbb{Z}_{(p)}}$ and $\operatorname{SL}_{2,\mathbb{Z}[\frac{1}{m}]}$ do not satisfy the ”complete reducibility” property. Over a field of $k$ characteristic zero, any finite dimensional comodule on $\operatorname{SL}(2,k)$ satisfies this property. In order to study torsion free comodules on $\operatorname{SL}_{2,\mathbb{Z}}$ we introduce the notion ”good filtration” and discuss the problem of classifying torsion free finite rank comodules with a good filtration. We give explicit non trivial examples of torsion free finite rank module with a good filtration, hence many nontrivial torsion free comodules has such a filtration, and the notion may therefore have general interest in the study of comodules on a reductive group scheme over a Dedekind domain. One may define the notion ”good filtration” for torsion free comodules on $\operatorname{SL}_{2,A}$ for any Dedekind domain $A$.